A spelunker is surveying a cave. She follows a passage 180 m straight west, then 210 m in a direction 450 east of south, and then 280 m at 300 east of north. After a fourth unmeasured displacement, she finds herself back where she started. Use a scale drawing to determine the magnitude and direction of the fourth displacement.


1
Expert's answer
2021-03-04T11:54:03-0500

Let's draw all the spelunker's displacements and find the unmeasured displacement:



Let's find the unknown "x" and "y" components of the unmeasured displacement of the spelunker:


"d_x=180\ m\cdot cos180^{\circ}+210\ m\cdot cos(360^{\circ}-45^{\circ})+280\ m\cdot cos30^{\circ}+x=0,""d_y=180\ m\cdot sin180^{\circ}+210\ m\cdot sin(360^{\circ}-45^{\circ})+280\ m\cdot sin30^{\circ}+y=0,""x=-211\ m, y=8.5\ m."

Then, the magnitude of the unmeasured displacement can be found from the Pythagorean theorem:


"d=\sqrt{d_x^2+d_y^2}=\sqrt{(-211\ m)^2+(8.5\ m)^2}=211.2\ m."

We can find the angle as follows:


"\theta=sin^{-1}(\dfrac{d_y}{d})=sin^{-1}(\dfrac{8.5\ m}{211.2\ m})=2.3^{\circ}."


The sign plus means that the unmeasured displacement has direction "2.3^{\circ} N\ of\ W".

Therefore, the unmeasured fourth displacement "d" has magnitude 211.2 m and direction of "2.3^{\circ} N\ of\ W".